OpenStudy (anonymous):

Can someone please help me solve this....D= A stone is thrown straight down from the edge of a roof, 725 ft above the ground, at a speed of 7 ft/sec. A. Given that the acceleration due to gravity is -32 ft/sec2, how high is the stone 4 seconds later? ft. B. At what time does the stone hit the ground? sec. C. What is the velocity of the stone when it hits the ground? ft/sec.

OpenStudy (anonymous):

write down all teh info u have is the best idea :) s= 725ft u= 7ft/s a=32/ft/s since that initial velocity is also acting downward, the gravity can also change to positive value. when t=4 \[s=ut + 1/2 at\]\[s=7ft/s+1/2(32ft/s²)(16s²)=263ft\] that is only the distance it travelled, actual height is : 725ft-263ft=462ft. to find its velocity after 4 second: \[v=u+at=7ft/s+32ft/s² \times 4s=135ft/s\] this velocity will be used as initial velocity in next calculation now: s=465ft u=135ft/s a=32ft/s² final velocity: \[v²=u² + 2as\]\[v=\sqrt{(135ft/s)²+2 \times 32ft/s² \times t}=218.616ft/s\] and with the final velocity,v we can use this equation: \[s=1/2(u+v)t\]\[462ft=1/2(135ft/s+218.161ft/s)t\]\[(462ft \times 2)/353.616ft/s=t=2.613\] 2.613 is the time fort the stone to reach the ground after 4 second, therefore, u need to plus 4 second back, and the total time for the whole free falling is 6.613s

OpenStudy (anonymous):

Hey, this is jst a correction on hei's answer from s = ut + 1/2 at2; then s = (7*4) +(1/2 *32*16) =28 +256 =284ft actual height = 725 - 284 = 441ft

OpenStudy (anonymous):

all the 263 need to change to 284 and 462 change to 441 2.613s change to 2.494s total time taken change to 6.494s, thanks for the correction :)